Solve by substitution.
x = y + 3 and 3x - 2y = 4

Answer by checkley71(8403)

X=Y+3 THEN 3(Y+3)-2Y=4 OR 3Y+9-2Y=4 OR Y=-5 THUS X=-5+3 OR X=-2